Grade 12Algebra

root of cosecx

shatharajupally vinay kumar
10 Years agoGrade 12

1 Answer

Anoopam Mishra
10 Years ago
\sqrt{cosec(x)} = \frac{1}{\sqrt{sin(x)}}